# PREPRINT VERSION

### v3

@@ Abstract
Marine surveys by robotic underwater and surface vehicles result in substantial quantities of coral reef imagery, however labeling these images is expensive and time-consuming for domain experts. Point label propagation is a technique that uses existing images labeled with sparse points to create augmented ground truth data, which can be used to train a semantic segmentation model. In this work, we show that recent advances in large foundation models facilitate the creation of augmented ground truth masks using only features extracted by the denoised version of the DINOv2 foundation model and K-Nearest Neighbors (KNN), without any pre-training. For images with extremely sparse labels, we present a labeling method based on human-in-the-loop principles, which greatly enhances annotation efficiency: in the case that there are 5 point labels per image, our human-in-the-loop method outperforms the prior state-of-the-art by 14.2% for pixel accuracy and 19.7% for mIoU; and by 8.9% and 18.3% if there are 10 point labels. When human-in-the-loop labeling is not available, using the denoised DINOv2 features with a KNN still improves on the prior state-of-the-art by 2.7% for pixel accuracy and 5.8% for mIoU (5 grid points). On the semantic segmentation task, we outperform the prior state-of-the-art by 8.8% for pixel accuracy and by 13.5% for mIoU when only 5 point labels are used for point label propagation. Additionally, we perform a comprehensive study into the impacts of the point label placement style and the number of points on the point label propagation quality, and make several recommendations for improving the efficiency of labeling images with points.
Index Terms:
Environmental Monitoring, Scene Understanding, Robotics, Human-in-the-Loop, Foundation Models, Semantic Segmentation
@@ I Introduction
This work establishes the utility of general foundation models for multi-species segmentation of domain-specific underwater imagery, with particular focus on label efficiency for the extremely sparse label setting (Fig. 1). Our contributions are summarized as follows:
@@
    We propose using a general-purpose foundation model to produce per-pixel deep features for coral images, demonstrating that these features are discriminative without training or fine-tuning on coral imagery. When combined with the basic K-Nearest Neighbors algorithm, these features are sufficient for creating accurate augmented ground truth masks, eliminating the need for custom superpixel algorithms.
@@
    We consider the extremely sparse point label setting, i.e. 5-25 points per image, and present a novel human-in-the-loop labeling regime, which integrates the expertise of the marine scientist with the model’s introspective uncertainty to identify informative locations for point labels. Our approach outperforms the prior state-of-the-art method by 14.2% pixel accuracy and 19.7% mIoU for the 5 point label setting, and by 8.9% and 18.3% when there are 10 points labeled in each image.
@@
    If human-in-the-loop labeling is not available, we find that leveraging the DINOv2 denoised features with a KNN results in improvements over the stage-of-the-art for propagation of small quantities of point labels per image. We see pixel accuracy improve by 2.7% and an improvement of 5.8% for mIoU (5 point labels per image); and 2.3% in pixel accuracy and 10.0% in mIoU (10 point labels per image) on UCSD Mosaics.
@@
    These improvements in point label propagation are reflected in the semantic segmentation task. A DeepLabv3+ model trained on augmented ground truth masks generated using DINOv2, KNN and our human-in-the-loop labeling regime significantly outperforms the prior works: pixel accuracy improves by 8.8% and mIoU improves by 13.5% (5 point labels per UCSD Mosaics image).
@@
    We conduct comprehensive experiments to assess the impact of the quantity and placement of point labels on the point propagation task, offering valuable recommendations for efficient annotation.
To foster future research in this area, we make our code publicly available [^2] https://github.com/sgraine/HIL-coral-segmentation.
@@ II-B Large Foundation Models
One such foundation model, DINOv2 [22], was trained on 142 million images, and is based on a Vision Transformer (ViT) model [44] trained with a discriminative self-supervised loss function [22]. This loss function combines an image-level objective function for features extracted from a student-teacher framework and a patch-level objective, in which patches from the input image are masked and the model must predict the missing image regions [22].
@@
While some works have explored the utility of foundation models for specialized problems [49, 47, 46], using DINOv2 for the task of underwater coral segmentation has not been investigated. As highlighted by Section II-A, coral imagery has unique and challenging visual characteristics, including detailed textures, poorly defined boundaries, and overlapping species, as seen in Fig. 2 [4, 20]. When applied to common imagery, such as on an example image from the Cityscapes [50] dataset, the DINOv2 model is able to effectively segment image, however when used directly on the coral images, the model is not able to generate useful segments (Fig. 2). This suggests that out-of-the-box usage of DINOv2 on this domain-specific task is not effective. However, our work investigates whether the feature representation of the foundation model DINOv2, trained on general images, can generate meaningful feature embeddings for coral images. We propose leveraging the deep embedding space of DINOv2 alongside sparse domain-expert point labels for point label propagation in coral imagery.
@@ III-A Method Overview
Figure 3: Schematic of Proposed Human-in-the-Loop Labeling Approach. We combine domain expert knowledge with the model’s internal uncertainty to improve point label selection. The process starts with inputting a coral image and having a marine scientist label up to 10 points centrally located on the largest instances (refer to Section V-B for an analysis of this value). A feature similarity map is then generated by computing cosine similarities between the labeled points and all other pixels. To promote exploration, we use a distance map in conjunction with the similarity map to create a combined probability mask for pixel selection. The chosen pixel is then sent back to the marine scientist for labeling, and the KNN is updated accordingly. Once the maximum number of points has been labeled, an augmented ground truth mask is created for use in training a semantic segmentation model.
@@ III-B Human-in-the-Loop Pixel Proposal
To implement this, we apply the method described in Section III-A to obtain, upsample, and normalize the per-pixel feature embeddings. We then calculate a cosine similarity map (Eq. 1) between the initially labeled pixels and every other pixel in the image. By inverting this map, we assign a higher probability of selection to pixel locations that exhibit low cosine similarity to the nearest labeled pixel:
@@ III-C Semantic Segmentation
After point label propagation (Stage One), we assess the augmented ground truth masks by training a fully supervised model for semantic segmentation (Stage Two). The model chosen for Stage Two can be tailored to meet the computational and inference time needs for deployment. In this study, we select the DeepLabv3+ model [56], a widely used architecture for semantic segmentation, and train it using the augmented ground truth masks as outlined in Section IV-A.
@@ IV-A2 Stage Two: Semantic Segmentation
We evaluate the propagated ground truth masks by training a model for semantic segmentation. We use TensorFlow to train the DeepLabv3+ model [56] and apply data augmentation techniques, including random horizontal and vertical flipping, and adjust the gain and gamma values within the range of $0.8-1.2$. The training process spans 500 epochs using the Adam optimizer with a learning rate of $0.001$. We report results on the test dataset based on the best epoch.
@@ IV-B Datasets
In this work, we use the UCSD Mosaics dataset, which contains multi-species coral images annotated with pixel-wise ground truth masks [51, 2]. In [2], the authors provide a version of the dataset where the large orthomosaics have been saved as smaller images, however we have identified some corrupted ground truth masks and have excluded them (this resulted in 219 being removed from the training set, resulting in 3,974 images, and 32 were removed from the test set, leaving 696 images; examples of the corruption can be seen in Fig. 4). We include a list of the corrupted filenames alongside our publicly available code [^5] https://github.com/sgraine/HIL-coral-segmentation. Each image measures 512 by 512 pixels and features 33 types of corals along with an class called ‘unknown’ or ‘unlabeled’. To maintain consistency with [2] and [4], we disregard this class during evaluation of the point label propagation task. We imitate the role of the domain expert in our human-in-the-loop labeling framework by obtaining the ground truth point label corresponding to the proposed location. For the first 10 pixels, we select the central pixel of the largest instances in the mask.
Refer to caption Refer to caption Refer to caption
Figure 4: Some of the ground truth masks in the UCSD Mosaics dataset showed signs of corruption, as illustrated at the top of these examples. The corresponding images and ground truth masks were removed from the training and test datasets.
@@ V Results
We compare our proposed approach to the point label propagation state-of-the-art in Section V-A and then Section V-B provides various ablation studies.
@@ V-A1 Stage One: Point Label Propagation
Our approach, which combines DINOv2 with our human-in-the-loop labeling regime, shows comparable computation times per image to the Point Label Aware Superpixel ensemble for small numbers of point labels (Table I). However, when applied to a large number of points i.e. 300 points, the computation time becomes excessively high due to the clustering in the deep feature space required to generate the feature similarity map at each iteration (Eq. 2). If there are a large number of point labels available, it is more suitable to use the grid-spaced version. We emphasize that the primary use case for human-in-the-loop labeling is to enhance performance in scenarios with very sparse point labels (5-25 points).
@@
Figure 6: A comparison is made between Fast MSS [3], Point Label Aware Superpixel (PLAS) [4], and our approach, which combines denoised DINOv2 features [23], K-Nearest Neighbors, and our Human-in-the-Loop labeling scheme. All methods are demonstrated on the same four examples. The top section shows point label propagation with 5 labels, while the bottom section displays propagation with 300 labels. Black circles indicate the point labels used for propagation within each output augmented mask. While all methods effectively propagate ground truth with 300 labels, the Fast MSS and PLAS methods struggle to produce usable augmented ground truth with only 5 labels. In contrast, our approach generates augmented ground truth that accurately resembles the ground truth even with the sparse 5 label setting.
@@ V-A2 Stage Two: Semantic Segmentation
TABLE II: Performance of Stage Two: Semantic Segmentation with DeepLabv3+ (Refer to Section IV-C for Metric Definitions), for 5 / 10 / 25 / 300 Point Labels. ‘F-MSS’ is Fast MSS [3], ‘PLAS’ is Point Label Aware Superpixels [4], and ‘D+NN’ is KNN with Denoised DINOv2 [23] (Ours).
Label PA mPA mIoU Method Style 5 / 10 / 25 / 300 5 / 10 / 25 / 300 5 / 10 / 25 / 300 F-MSS Grid 48.05 / 51.06 / 57.41 / 85.34 5.97 / 10.39 / 14.83 / 63.54 4.41 / 8.56 / 13.04 / 52.16 PLAS - Ens. Grid 65.73 / 70.18 / 73.60 / 83.72 27.96 / 37.64 / 48.53 / 66.89 19.48 / 26.01 / 36.27 / 52.83 D+NN (Ours) Grid 72.24 / 77.64 / 85.93 / 85.93 32.16 / 41.61 / 52.59 / 62.99 25.66 / 34.80 / 43.41 / 54.07 D+NN (Ours) HIL 74.53 / 71.04 / 81.69 / 86.29 41.47 / 51.62 / 54.31 / 63.47 32.96 / 38.21 / 45.46 / 54.62

### v4

# Human-in-the-Loop Segmentation of Multi-species Coral Imagery

@@ Abstract
Marine surveys by robotic underwater and surface vehicles result in substantial quantities of coral reef imagery, however labeling these images is expensive and time-consuming for domain experts. Point label propagation is a technique that uses existing images labeled with sparse points to create augmented ground truth data, which can be used to train a semantic segmentation model. In this work, we show that recent advances in large foundation models facilitate the creation of augmented ground truth masks using only features extracted by the denoised version of the DINOv2 foundation model and K-Nearest Neighbors (KNN), without any pre-training. For images with extremely sparse labels, we use human-in-the-loop principles to enhance annotation efficiency: if there are 5 point labels per image, our method outperforms the prior state-of-the-art by 19.7% for mIoU. When human-in-the-loop labeling is not available, using the denoised DINOv2 features with a KNN still improves on the prior state-of-the-art by 5.8% for mIoU (5 grid points). On the semantic segmentation task, we outperform the prior state-of-the-art by 13.5% for mIoU when only 5 point labels are used for point label propagation. Additionally, we perform a comprehensive study into the number and placement of point labels, and make several recommendations for improving the efficiency of labeling images with points.
@@ I Introduction
This work establishes new methodological approaches for foundation model adaptation in specialized domains, demonstrating how general representations can be effectively combined with domain-specific expert knowledge for efficient annotation. For the first time, we address the problem of label efficiency in the extremely sparse label setting (Fig. 1) through principled human-machine collaboration. Our contributions are summarized as follows:
@@
    Foundation Model Adaptation: We propose using a general-purpose foundation model to produce per-pixel deep features for coral images, demonstrating that these features are discriminative without training or fine-tuning on coral imagery. When combined with the basic K-Nearest Neighbors algorithm, these features are sufficient for creating accurate augmented ground truth masks, eliminating the need for custom superpixel algorithms.
@@
    Human-in-the-Loop Framework: We present a novel human-in-the-loop labeling regime for the extremely sparse point label setting, i.e. 5-25 points per image, which integrates the expertise of the marine scientist with the model’s introspective uncertainty to identify informative locations for point labels. Our approach outperforms the prior state-of-the-art method by 14.2% pixel accuracy and 19.7% mIoU for the 5 point label setting, and by 8.9% and 18.3% when there are 10 points labeled in each image.
@@
    Robust Performance without Human Guidance: If human-in-the-loop labeling is not available, we find that leveraging the DINOv2 denoised features [22, 23] with a KNN results in improvements over the stage-of-the-art for propagation of small quantities of point labels per image. We see pixel accuracy improve by 2.7% and an improvement of 5.8% for mIoU (5 point labels per image); and 2.3% in pixel accuracy and 10.0% in mIoU (10 point labels per image) on UCSD Mosaics [24, 2].
@@
    End-to-End Semantic Segmentation: These improvements in point label propagation are reflected in the semantic segmentation task. A DeepLabv3+ model [25] trained on augmented ground truth masks generated using DINOv2 [22, 23], KNN and our human-in-the-loop labeling regime significantly outperforms the prior works: pixel accuracy improves by 8.8% and mIoU improves by 13.5% (5 point labels per UCSD Mosaics [24, 2] image).
@@
    Comprehensive Analysis: We conduct comprehensive experiments to assess the impact of the quantity and placement of point labels on the point propagation task, offering valuable recommendations for efficient annotation.
An earlier version of this work was presented as a conference paper [26]. The major contributions compared to the conference version are:
  1. 1.
    Extension of the proposed algorithm to a complete two-stage semantic segmentation pipeline and thorough analysis of the second stage:
    • Addition of Stage Two enables inference on unlabeled images at test time
    • Implementation of a semantic segmentation model for automated analysis of coral imagery (described in Section III-C and Section IV-A2)
    • Integration of both stages into a cohesive framework for practical application (shown in Fig. 5)
  2. 2.
    Comprehensive experimental evaluation and analysis:
    • Doubled the quantitative and qualitative results presented (Tables II and IV, and expansion of Figs. 9 and 10)
    • Evaluation of various semantic segmentation architectures and their backbones (Section V-A2 and Section V-B6)
    • Comparative analysis of inference speeds for practical deployment considerations (Table IV)
  3. 3.
    Expanded ablation studies for the point label propagation approach:
    • New analysis of clustering parameters ($k$ values) across different quantities of point labels (Fig. 13)
    • Investigation of minimum number of human-labeled initial points required (Fig. 14)
    • Extended evaluation of feature extractors, including DINOv2 variants with Registers [22, 27] (Fig. 10)
  4. 4.
    Enhanced visual results and in-depth discussion:
    • New demonstrations of foundation model limitations on complex coral imagery (Figs. 2 and 3 provide analysis of Segment Anything 2 [28, 29] and DINOv2 [22] results when used out-of-the-box)
    • Analysis of positional embedding artifacts in the feature space (Fig. 11)
    • Additional examples illustrating dataset cleaning (Fig. 7) and visualization of ground truth masks (Fig. 6)
    • Additional discussion of impact, limitations and future work for the proposed method (Section VI).
To foster future research in this area, we make our code publicly available: https://github.com/sgraine/HIL-coral-segmentation.
@@ II-A Semantic Segmentation of Coral Imagery
There have been recent advances in large, general instance segmentation models e.g. the Segment Anything (SAM) [28] and Segment Anything 2 [29] models. It is important to highlight the difference between the instance segmentation and semantic segmentation tasks. Instance segmentation groups pixels in an image into distinct segments without classifying the object. Semantic segmentation instead predicts the presence of specific classes of interest, and localizes those objects by denoting the pixels which belong to each class. We evaluate the performance of SAM 2 on coral imagery in Fig. 2, which demonstrates that the output of these models does not provide class information, and that the segments produced do not align well with the intended ground truth masks.
Figure 2: Instance segmentation by Segment Anything 2 [29] on UCSD Mosaics [24, 2] coral images. This figure shows that Segment Anything 2 does not produce usable segments in these images, and also does not provide any class predictions for the segments. Refer to Fig. 6 for the color legend for the ground truth masks.
@@ II-B Large Foundation Models
One such foundation model, DINOv2 [22], was trained on 142 million images, and is based on a Vision Transformer (ViT) model [49] trained with a discriminative self-supervised loss function [22]. This loss function combines an image-level objective function for features extracted from a student-teacher framework and a patch-level objective, in which patches from the input image are masked and the model must predict the missing image regions [22]. There have been extensions and improvements of the DINOv2 architecture: the Denoising Vision Transformers approach [23] removes positional artefacts from deep features, and can be applied on top of any ViT; and Darcet et al. [27] propose using registers, which are additional tokens the model can use to store global information, therefore reducing artefacts in ViT features.
@@
While some works have explored the utility of foundation models for specialized problems [54, 52, 51], using DINOv2 for the task of underwater coral segmentation has not been investigated. As highlighted by Section II-A, coral imagery has unique and challenging visual characteristics, including detailed textures, poorly defined boundaries, and overlapping species, as seen in Fig. 3 [4, 20]. When applied to common imagery, such as on an example image from the Cityscapes [55] dataset, the DINOv2 model is able to effectively segment image, however when used directly on the coral images, the model is not able to generate useful segments (Fig. 3). This suggests that out-of-the-box usage of DINOv2 on this domain-specific task is not effective. However, our work investigates whether the feature representation of the foundation model DINOv2, trained on general images, can generate meaningful feature embeddings for coral images. We propose leveraging the deep embedding space of DINOv2 alongside sparse domain expert point labels for point label propagation in coral imagery.
@@ III-A Method Overview
Figure 4: Schematic of Proposed Human-in-the-Loop Labeling Approach. We combine domain expert knowledge with the model’s internal uncertainty to improve point label selection. The process starts with inputting a coral image and having a marine scientist label up to 10 points centrally located on the largest instances (refer to Section V-B for an analysis of this value). A feature similarity map is then generated by computing cosine similarities between the labeled points and all other pixels. To promote exploration, we use a distance map in conjunction with the similarity map to create a combined probability mask for pixel selection. The chosen pixel is then sent back to the marine scientist for labeling, and the KNN is updated accordingly. Once the maximum number of points has been labeled, an augmented ground truth mask is created for use in training a semantic segmentation model. The semantic segmentation architecture is trained end-to-end on pairs of images and augmented ground truth masks (as seen in Fig. 5, which demonstrates that this model is then used to perform inference on new images).
@@ III-B Human-in-the-Loop Pixel Proposal
To implement this, we apply the method described in Section III-A to obtain, upsample, and normalize the per-pixel feature embeddings. We then calculate a cosine similarity map (1) between the initially labeled pixels and every other pixel in the image. By inverting this map, we assign a higher probability of selection to pixel locations that exhibit low cosine similarity to the nearest labeled pixel:
@@ III-C Semantic Segmentation
After point label propagation (Stage One), we use the augmented ground truth masks to train a fully supervised model for semantic segmentation (Stage Two). The model chosen for Stage Two can be tailored to meet the computational and inference time needs for deployment. We evaluate the performance of various semantic segmentation architectures in Section V-B. We compare the DeepLabv3+ architecture [60] with ResNet50 [61], Inceptionv3 [62] and Mobilenetv2 backbones with UNet [63] and LinkNet [64] with ResNet50 [61] backbone. We train the semantic segmentation architectures using the augmented ground truth masks, as outlined in Section IV-A.
Figure 5: Schematic of Full Semantic Segmentation Framework. We combine the human-in-the-loop point label propagation approach (Stage One) with a semantic segmentation architecture (Stage Two). The semantic segmentation architecture is trained on the augmented ground truth masks to enable inference on unlabeled imagery. Refer to Section V-B for an analysis of various semantic segmentation architectures.
@@ IV-A2 Stage Two: Semantic Segmentation
We evaluate the propagated ground truth masks by training a model for semantic segmentation. We use TensorFlow to train a DeepLabv3+ model [60] with ResNet50 backbone [61]. We also compare alternate semantic segmentation architectures and backbones in the ablation study in Section V-B, and we report the inference times of these architectures in frames per second (fps, a higher number indicates faster inference), when inference is performed on a NVIDIA A100 GPU. During training we apply data augmentation techniques, including random horizontal and vertical flipping, and adjust the gain and gamma values within the range of 0.81.20.8-1.2. The training process spans 500 epochs using the Adam optimizer with a learning rate of $0.001$. We report results on the test dataset based on the best epoch.
@@ IV-B Datasets
Recent works have contributed datasets of underwater imagery for semantic segmentation [33, 38, 39], however these datasets do not include multiple species of corals. As coral imagery is highly complex, with overlapping instances and poorly defined outlines, it is extremely time-consuming and expensive for marine scientists to provide pixel-wise dense annotations to accompany the training images. Therefore, the vast majority of multi-species coral datasets are labeled only with sparse point labels e.g. [67, 58]. The UCSD Mosaics dataset is the only multi-species coral dataset with accompanying pixel-wise ground truth segmentation masks [2, 4]. We acknowledge that this is a significant limitation and challenge for the field of multi-species segmentation of coral imagery, and there is a clear need for more multi-species, densely labeled datasets of coral imagery fore comparison of approaches. In this work, we therefore use the UCSD Mosaics dataset [24, 2] for evaluation of our proposed approach and comparison to prior works. The ground truth masks are needed to accurately evaluate the performance of semantic segmentation models. The colorised legend for the pixel-wise ground truth segmentation masks is in Fig. 6 – we refer to this legend throughout this paper when showing results on this dataset.
Figure 6: Legend for the UCSD Mosaics dataset [24, 2]. This table displays the color coding for the semantic classes referenced throughout the paper. The legend corresponds to the classes present in the ground truth masks, point label propagation and segmentation outputs shown in Figs. 1-4, 7 and 9-10. Only classes appearing in the visualizations included in this paper are shown.
In [2], the authors provide a version of the dataset where the large orthomosaics have been saved as smaller images, however we have identified some corrupted ground truth masks and have excluded them (this resulted in 219 being removed from the training set, resulting in 3,974 images, and 32 were removed from the test set, leaving 696 images; examples of the corruption can be seen in Fig. 7). We include a list of the corrupted filenames alongside our publicly available code [^4] https://github.com/sgraine/HIL-coral-segmentation. Each image measures 512 by 512 pixels and features 33 types of corals along with an class called ‘Unidentified’. To maintain consistency with [2] and [4], we disregard this class during evaluation of the point label propagation task. We imitate the role of the domain expert in our human-in-the-loop labeling framework by obtaining the ground truth point label corresponding to the proposed location. For the first 10 pixels, we select the central pixel of the largest instances in the mask.
@@ IV-C Evaluation Metrics
Refer to caption Refer to caption Refer to caption
Figure 7: Some of the ground truth masks in the UCSD Mosaics dataset showed signs of corruption, as illustrated at the top of these examples. The corresponding images and ground truth masks were removed from the training and test datasets. Refer to Fig. 6 for the color legend.
@@ V Results
We compare our proposed approach to the point label propagation state-of-the-art in Section V-A and then Section V-B provides various ablation studies. We describe our findings on the effect of point label quantity in Section V-C and the effect of point label placement style in Section V-D.
@@ V-A1 Stage One: Point Label Propagation
Our approach, which combines DINOv2 with our human-in-the-loop labeling regime, shows comparable computation times per image to the Point Label Aware Superpixel ensemble for small numbers of point labels (Table I). However, when applied to a large number of points i.e. 300 points, the computation time becomes excessively high due to the clustering in the deep feature space required to generate the feature similarity map at each iteration (2). If there are a large number of point labels available, it is more suitable to use the grid-spaced version. We emphasize that the primary use case for human-in-the-loop labeling is to enhance performance in scenarios with very sparse point labels (5-25 points).
@@
Figure 9: A comparison is made between Fast MSS [3], Point Label Aware Superpixel (PLAS) [4], and our approach, which combines denoised DINOv2 features [23], K-Nearest Neighbors, and our Human-in-the-Loop labeling scheme. All methods are demonstrated on the same four examples. The top section shows point label propagation with 5 labels, while the bottom section displays propagation with 300 labels. Black circles indicate the point labels used for propagation within each output augmented mask. While all methods effectively propagate ground truth with 300 labels, the Fast MSS and PLAS methods struggle to produce usable augmented ground truth with only 5 labels. In contrast, our approach generates augmented ground truth that accurately resembles the ground truth even with the sparse 5 label setting. Refer to Fig. 6 for the color legend.
@@ V-A2 Stage Two: Semantic Segmentation
TABLE II: Performance of Stage Two: Semantic Segmentation with DeepLabv3+ (Refer to Section IV-C for Metric Definitions), for 5 / 10 / 25 / 300 Point Labels. ‘F-MSS’ is Fast MSS [3], ‘PLAS’ is Point Label Aware Superpixels [4], and ‘D+NN’ is KNN with Denoised DINOv2 [23] (Ours). For all rows, the semantic segmentation architecture used is the Deeplabv3+ architecture [60] with ResNet50 backbone [61].
Label PA mPA mIoU Method Style 5 / 10 / 25 / 300 5 / 10 / 25 / 300 5 / 10 / 25 / 300 F-MSS [3] Grid 48.05 / 51.06 / 57.41 / 85.34 5.97 / 10.39 / 14.83 / 63.54 4.41 / 8.56 / 13.04 / 52.16 PLAS - Ens. [4] Grid 65.73 / 70.18 / 73.60 / 83.72 27.96 / 37.64 / 48.53 / 66.89 19.48 / 26.01 / 36.27 / 52.83 D+NN (Ours) Grid 72.24 / 77.64 / 85.93 / 85.93 32.16 / 41.61 / 52.59 / 62.99 25.66 / 34.80 / 43.41 / 54.07 D+NN (Ours) HIL 74.53 / 71.04 / 81.69 / 86.29 41.47 / 51.62 / 54.31 / 63.47 32.96 / 38.21 / 45.46 / 54.62